OpenStudy (anonymous):

For the reaction C4H10 <-> C4H6 + 2H2 the equilibrium constant Kc is 1.00x10^-6 at 600ºC. If 1.00 mol of C4H6 and 1.00 mol of H2 were introduced into a 1.00L flask at 600ºC, what would be the equilibrium concentration of each gas?

OpenStudy (anonymous):

I do it with RICE usually...So: R) C4H10 <-> C4H6 + 2H2 I) ----- 1M 1M C) E) I don't know how to get the C and E without the Initial Molarity of C4H10

OpenStudy (anonymous):

1.00 x 10^-6 = (1+x)(1+2x)^2/-x Can you help me from here? @hartnn

hartnn (hartnn):

you sure thats correct equation ?? because you will get a really weird answer for thsi

OpenStudy (anonymous):

Pretty sure... R) C4H10 <-> C4H6 + 2H2 I) ----- 1M 1M C) -x +x +2x E) -x 1+x 1+2x Kc = [C4H6][H2]^2/[C4H10] @ash2326 wanna try n help me if u know if this is right? <3

hartnn (hartnn):

this can become easier, if we approximate things, like \(\large 10^{-6} =0\) approximately, compared to 1

hartnn (hartnn):

so we would only have \((1+x)(1+2x)^2 =0\) which is very easy to solve

hartnn (hartnn):

\(1+x = 0, \quad x =...\\ 1+2x = 0, \quad x=....\)